# Improving sampling of binding free energy differences between covalently bound ligands in alternate binding pockets using MT-REXEE

### v1

@@ 3.1 System Set-up
We used two systems to test this approach, MUP1-thiazole and FabB-acyl complexes described in the introduction. The MUP1-thiazole complex features a cyclic ligand bound to a relatively rigid $\beta$-barrel protein domain for which convergence of free estimates requires easily accessible computational resources 28. We used the FabB-acyl complex allows us to demonstrate the conformational sampling enhancement of the MT-REXEE approach as it features high free energy barriers for acyl chain rearrangement, which as described above are difficult to sample with existing methods. The protein-thiazole complexes were derived from the crystal structure (PDB 1IO6). The FabB covalent intermediate complexes were derived from the crystal structure of WT FabB with a covalently bound C12 acyl chain (PDB 1EK4). MODELLER was then used to apply the G107M mutation. All small molecules were parameterized using GAFF2 29 and the protein was parameterized using the Amber ff14SB force field30. The relative binding free energy topologies for all systems were created using PMX 31 followed by post-processing to remove all dihedral energies between dummy and real atoms within the hybrid topologies. The treatment of dummy atoms in this study follows the practices presented in earlier works to minimize any effects on the partition function that are not canceled out by alchemical transformations in different environments  32, 33. The unit cell boxes for all simulations conducted in solvent were constructed to maintain a minimum distance of 1 nm between the molecule and the periodic boundary and then filled with TIP3P 34 solvent. All simulations were run in a modified version of GROMACS 2022.5 35, which can be found as a branch on the official GROMACS GitLab repository; all code changes in this branch have been merged into the main repository as of GROMACS 2025.0. This modification allows the acceptance of the MDP parameter init_histogram_counts, which is crucial for performing Wang-Landau weight equilibration across iterative trajectories.
@@
We initiated the FabB covalent intermediate complex with the acyl chain in two different, non-overlapping pockets, which are referred to as pocket A and pocket B (Figure 3A). Equilibration was performed for the FabB–acyl complex in pocket A with 50 ns using standard MD at $\lambda=0$. From this simulation, three random frames were selected in the first, second, and last third of the trajectory as the initial configuration for each of the three replicas. We then run MT-REXEE simulations for the FabB–acyl complex in pocket A. These simulations sample both pocket A and pocket B. We extract one frame in which each complex samples pocket B from these MT-REXEE simulations. We then ran 50 ns of standard MD on each of these pocket B configurations at $\lambda=0$ and again extracted a random frame from the first, second, and last third of the MD trajectories as the initial configurations.
@@ 3.2 Free Energy Protocol
We performed both EE and MT-REXEE simulations in triplicate for each system. The EE settings used were identical for both the MT-REXEE and independent EE simulations with a Wang-Landau scale of 0.82 and a Wang-Landau ratio of 0.85 for the equilibration for the FabB system, and 0.8 was used for both settings for the MUP1 inhibitor binding free energy. These parameters were tuned to ensure relatively even sampling between $\lambda$ intermediate states following weight equilibration. Slower weight equilibration through the increasing of Wang-Landau scale and ratio parameters extends equilibration times but produces more even $\lambda$ state sampling once equilibrated weights are obtained. The weight equilibration phase was ended when the Wang-Landau $\delta$ was less than 0.001 for both systems, a cutoff determined from previous experience. The MUP1 inhibitors were run for 30 ns and 50 ns after equilibration for the solvent and protein complex simulations, respectively. We simulated the FabB complexes in the covalent intermediate state, but since the ligand was previously covalently bound to the protein ACP, the“solvent” simulation state was defined as ACP with the acyl chain in solvent. The FabB systems were run for 50 ns and 100 ns for the solvent and enzyme-bound states, respectively. All free energy estimates were calculated using the MBAR implementation in alchemlyb 38.
The primary feature added to the MT-REXEE package in this study is the use of a new swapping criterion between simulations. In the previously developed exhaustive approach, swaps are only attempted for the last frame of each iteration. This new swapping criterion is named random range (RR), and rather than only allowing swaps between trajectories in the last frame of an iteration, we instead allow swaps at random compatible frames (i.e. endpoints of two neighboring simulations) in a defined portion of the trajectory. We used the last 50% of the trajectory to avoid swaps being highly correlated to the previous iteration. Regardless of the location of the trajectory used for swapping, all data from each iteration is used to compute free energy estimates and any additional analysis; only trajectory swapping and continuation use the randomly sampled exchange-compatible frame. For the MUP1 system, we used redundant end states (5 intermediate $\lambda$ states as well as 5 states each at $\lambda$=0 and $\lambda$=1) in combination with the exhaustive swapping method as we did in our previous MT-REXEE paper 5 to compare to the random range swapping approach without redundant end states. For the FabB–acyl systems, we only ran MT-REXEE simulations with the RR method and with single $\lambda$ end-states.
@@ 3.3 Configurational Analysis
X-ray crystallography structures of the FabB–acyl-ACP complex have previously identified two distinct binding tunnels for the acyl chain. We predict that these distinct binding tunnels are also occupied in the covalent intermediate states, but there are no experimental structures defining this alternate binding pocket, which we term pocket B, in the covalent intermediate state. We used the MT-REXEE simulations with the acyl chain initialized in the canonical binding pocket, which we term pocket A. We then visually labeled a subset of trajectory frames (12–15) for all chain lengths as being in either pocket A or B in order to then determine a quantitative measurement. We determined that the distance between carbon 4 in the acyl chain and residue THR300 could be used to differentiate the binding pockets (Figure 8A). We evaluated several distance thresholds between pairs of atoms in the acyl chain and FabB as well as acyl internal dihedrals as metrics for pocket determination, but the distance between C4 and THR300 metric was determined to have the least overlap in values between pocket A and pocket B. We plotted the distance distribution of these residues in EE simulations which were visually verified to have all frames within either pocket A or B, as well as for frames from our MT-REXEE simulations which sample both pockets to determine the threshold of 0.85 nm (Figure 8B). Several conformations for longer chains (8 or more carbons) were identified as belonging in pocket A or B by this criterion, but did not match existing known structures for the pockets. Thus, we also applied a chain compaction criterion that the distance between carbon 2 and 8 should be shorter than 0.6 nm in pocket B and longer than 0.75 nm in pocket A determined from available crystal structures. Any trajectory frames in which this distance to THR300 is greater than 0.85 nm and the compaction distance is greater than 0.75 nm are considered in pocket A. Any frames with a distance to THR300 less than 0.85 nm and a compaction distance of less than 0.6 nm is considered in pocket B. Any frames for which neither pocket criteria were satisfied were discarded from free energy analysis, as they are in neither pocket (<<2% of trajectories).
@@ 4 Results
4.1 Random range swapping improves round trip time and statistical convergence compared to previous exchange methods
@@
In this section, we compare the results for running binding free energy calculations for a set of small molecule ligands with the protein MUP1 using both the previous introduced exhaustive method and the random range (RR) method (Figure 2A–B). We ran this comparison to determine any changes in performance and ensure that this new algorithm for selecting swappable states does not introduce additional error. The RR method no longer requires redundant end states. This eliminates the increase in Wang-Landau weight equilibration times observed in our earlier work 5. The result of this improvement is that MT-REXEE with the RR exchange method (MT-REXEE RR) has equivalent or lower total Wang-Landau weight equilibration times compared to independent EE simulations (Figure 9). We note that weight equilibration times are calculated using the full iteration in which the weights reach the equilibration criteria, even if conformational exchanges occur at a frame earlier than the final frame of the iteration. Only the E-F transformation had a potentially statistically significant difference between the weight equilibration times using EE and MT-REXEE RR ($p=0.083$). We hypothesize that the decrease in weight equilibration times for the E-F transformation is likely due to the improved conformational space sampling using RR. This transformation is the most chemically distinct of all the MUP1 ligands evaluated in this study and for this reason, benefits most from the increased conformational sampling.
@@
Figure 2: Random range swapping yields statistically equivalent free energies while significantly decreasing the round trip time through alchemical states. (A) A diagram of the sequence of ligand transformations used in binding free energy estimates with the protein MUP1 to validate the random range swapping scheme. The curved arrows show where MT-REXEE swaps occur. (B) The structure of MUP1 in complex with ligand A. (C) We compare the binding free energy estimates using $E$, MT-REXEE with the exhaustive swapping scheme, and MT-REXEE with the RR swapping scheme. We observe no significant differences between the free energy estimates for all three methods as the values are all within statistical error of one another. (D) We observe a significant (p<<0.001) decrease in the round-trip times from 10-20 ns to 3-5 ns for both solvent and MUP1 complex simulations using the RR swapping scheme compared to the previous exhaustive scheme.
@@
The binding free energy estimates are statistically equivalent for EE and MT-REXEE with both exhaustive and RR swapping methods, but various convergence metrics indicate that less simulation time is necessary for equal accuracy with RR (Figure 2C). We measured the convergence of these free energy estimates both along the length of individual trajectories as well as convergence between trajectories initiated with significantly different initial configurations. Importantly, we observe that for both solvent and, even more significantly for the protein complex simulations, the free energy estimates converge significantly more quickly, resulting in closer estimates between the estimates for the first and last half of the trajectories for MT-REXEE RR compared to both EE and the MT-REXEE simulations using the exhaustive method (Figure 10A–B). We also notice a significant decrease in the standard error of the estimate between replicates of between 40 and 160% compared to EE (Figure 10C). These independent replicates were initiated with the same initial configurations extracted from standard MD trajectories separated by a minimum of 20 ns. The exchanges between distinct alchemical transformations in the MT-REXEE method enable increased sampling of the conformational space, which produces faster convergence. The standard error is also reduced by between 16-90% when using the exhaustive swapping method with redundant end states compared to EE. This is likely because the increase in conformational sampling is balanced by the decreased sampling of intermediate $\lambda$ states due to the addition of redundant end states. This results in decreased conformational sampling than MT-REXEE RR, but still faster conformational sampling than EE. This is consistent with the observed significant decrease in the round-trip times for the MT-REXEE RR method compared to the exhaustive swapping method (Figure 2D). We expect the over 4-fold decrease in round-trip time is primarily responsible for the faster increased conformational sampling in MT-REXEE RR compared to the MT-REXEE exhaustive method.
@@ 4.2 Enhanced Sampling of FabB with MT-REXEE
We apply the MT-REXEE method to an example system in which there is a known sampling limitation. Prior evidence suggests that in the FabB-acyl-ACP complex the acyl chain samples two distinct binding tunnels, pocket A and pocket B. Here we examine the FabB-acyl intermediate complex, after the transfer of the acyl chain from ACP to FabB through the formation of a covalent bond with CYS163 on FabB (Figure 3A). Testing sampling with the FabB-acyl complex simplifies assessing the sampling compared to the FabB-acyl-ACP complex, as the motions of ACP produce additional sampling issues as the FabB–ACP interface is large and flexible. Using the FabB-acyl complex allows us to focus on only the motions of the acyl chain. The FabB-acyl complex is also the well-known intermediate conformation before the condensation reaction with malonyl-ACP and the covalent intermediate to extend the acyl chain by two carbons, making it a biologically important state 10.
@@
We did not know whether the FabB-acyl covalent intermediate step would sample pocket B as was observed with the FabB–acyl-ACP complex, but previous mutational data did support sufficiently stable binding to enable production of acyl chains up to 16 carbons in length via FabB with a mutation which blocked pocket A through steric hindrance. We verified that the transition between the pocket A and B is too slow to sample with direct simulation using standard MD alone within reasonable simulation lengths (500 ns; Figure 3B). Therefore, this study uses MT-REXEE with random range swapping to evaluate the relative stability of 2, 4, 6, 8, 10, 12, 14, and 16 carbon length acyl chain substrates covalently bound to FabB. We only attach one substrate in a given simulation even though FabB is simulated in its native dimer conformation. This is motivated by existing evidence which suggests there may be negative cooperation between the two dimer members and it is thus more likely that only one reacts at a time.
@@
We performed the 500 ns standard MD simulations for FabB-acyl complexes with acyl substrates from 2–16 carbons in length, seven simulations in all. All simulations were initiated with the same initial configuration as the corresponding MT-REXEE simulation in pocket A, with the removal of dummy atoms required for MT-REXEE. We ran three replicates with distinct initial configurations for all chain lengths (see Methods). The corresponding MT-REXEE simulations were run for only 100 ns (in parallel) for each transformation of chain length. We then analyzed the trajectories to determine the occupancy of pocket A, B, or neither for each chain length substrate. For MT-REXEE simulations, we specifically compared the pocket occupancy for the end states and not for intermediate $\lambda$ states, as these are irrelevant to the physical occupancies.
@@
Figure 3: (A) The binding pocket of the FabB–acyl covalent intermediate is shown with a 12 carbon acyl substrate in both pocket A (orange) and pocket B (brown) in the context of the overall protein context. The structures shown are the centroids from MT-REXEE simulations of the C10-12 transformation for $\lambda=1$. (B) Occupancy of pocket B as a function of substrate chain length using both standard MD simulations and MT-REXEE enhanced sampling as described in section $3.$ Standard MD is not able to access pocket B within 500 ns for any substrate longer than 4 carbons, but we readily see sampling in pocket B for all chain lengths within only 100 ns of MT-REXEE simulation. (C) A close-up image of the binding pocket in (A) highlighting several residues on both the catalytically active monomer and adjacent monomer that restrict the acyl chain’s conformational flexibility and prevent interconversion during MD simulations of at least 500 ns.
@@
MT-REXEE and standard MD show statistically equivalent sampling of pocket B only for the 2 carbon acyl chain substrate (Figure 3B). We found there is minimal sampling of pocket B for the 4 carbon substrate in standard MD and for all longer substrates there is no sampling of pocket B in standard MD started from pocket A. MT-REXEE thus enables significant sampling in pocket B for all chain lengths studied while utilizing only 20% of the trajectory length as the standard MD simulations.
@@
We hypothesize that the inability of standard MD to sample both pockets is due to the presence of several acyl-chain-stabilizing interactions present for all acyl chain lengths 4 carbons or longer (Figure 3C). For both MD and MT-REXEE simulations, the 2-carbon acyl chain substrates are readily able to sample both pockets A and B with relatively high frequency. The presence of the conformational swaps between the 2 carbon state and longer chain substrates allows for conformational sampling of both pockets at all chain lengths. MT-REXEE enables simulation trajectories to sample pocket B when simulations are initiated in the canonical binding pocket A, thus requiring no prior knowledge of the presence or location of the alternate pocket. In contrast, existing knowledge of both the location of pocket B and a reasonable transition path would be necessary for employing a configurational variable dependent method. The lack of directly applied configurational bias along a known CV, like that used in metadynamics, also provides increased confidence that these alternate conformations are not artifacts of the enhanced sampling method itself.
@@
In MT-REXEE simulations, switching between pockets is primarily limited by the round-trip time between alchemical simulations of different chain lengths, as we can only generally only switch between pockets in simulations of the shortest chain. Additionally, we want to emphasize that the success of the MT-REXEE method is heavily correlated to the suitability of the applied equilibrated EE weights. An increase in round-trip time correlates with an increased imbalance between $\lambda$ state sampling within the simulations (Figure 11). The imbalance in $\lambda$ state sampling can be minimized by adjusting Wang-Landau equilibration parameters. Optimization of these parameters is system-specific. Further optimization of these parameters would likely further minimize this effect, but would also increase weight equilibration times. This necessitates a balance between the efficiency of weight equilibration and the accuracy of equilibrated weights.

### v2

@@ 3.1 System Set-up
We used two systems to test our approaches: (i) the complex formed between mouse major urinary protein 1 (MUP1) and thiazole-based pheromones which we previously studied with MT-REXEE14 and (ii) the FabB-acyl complexes described in the introduction. The MUP1-thiazole complex contains a cyclic ligand bound to a rigid $\beta$-barrel protein domain for which convergence of free energy estimates requires easily accessible computational resources 34. The FabB-acyl complex is an ideal example to demonstrate the conformational sampling enhancement of the MT-REXEE approach. This system features high free energy barriers for acyl chain rearrangement, which create trapped conformational states for which the transitions are prohibitively difficult to sample with existing methods. We derived the MUP1 complexes from the crystal structure this protein bound to 2-sec-butyl-4,5-dihydrothiazole (PDB 1I06). We derived the FabB-acyl complexes from the crystal structure of FabB covalently bound to a C12 acyl chain (PDB 1EK4). For FabB, we used MODELLER to apply the G107M mutation and we used Avogadro to shrink or elongate the acyl chain to generate C4-16 initial configurations. All small molecules were parameterized using GAFF2 17 and the protein was parameterized using the Amber ff14SB force field23. The relative binding free energy topologies for all systems were created using PMX 15 followed by post-processing to remove all dihedral energies between dummy and real atoms within the hybrid topologies. The treatment of dummy atoms in this study follows the practices presented in earlier works to minimize any effects on the partition function that are not canceled out by alchemical transformations in different environments 13, 27. The unit cell boxes for all simulations conducted in solvent were constructed to maintain a minimum distance of 1 nm between the molecule and the periodic boundary and then filled with TIP3P 21 solvent. All simulations were run in a modified version of GROMACS 2022.5 4, which can be found as a branch on the official GROMACS GitLab repository. This modification allows the acceptance of the MDP parameter init_histogram_counts, which is crucial for performing Wang-Landau weight equilibration across iterative trajectories. Importantly, all code changes in this branch have been merged into the main code base as of GROMACS 2025.0. The MT-REXEE implementation can be found within the ensemble_md GitHub package github.com/shirtsgroup/ensemble_md.
@@
We initiated the FabB-acyl complex with the acyl chain in two different, non-overlapping pockets, which are referred to as pocket A and pocket B (Figure 3A). An initial standard MD simulation was performed for the FabB-acyl complex in pocket A with 50 ns at $\lambda=0$. The purpose of this simulation was to extract decorrelated frames from the MD trajectory to use as initial configurations for our MT-REXEE simulations. A duration of 50 ns was selected to increase the probability of sampling alternate metastable states. The mean transition time between pockets A and B for the 2 carbon chains was approximately $11\pm 5$ ns, which gives us an approximation for the timescale of significant acyl chain rearrangements. We sought to extract three frames separated by 11-16 ns of MD simulation, requiring a total of 48 ns, which was rounded to 50 ns for simplicity. From this simulation, three random frames were selected in the first, second, and last third of the trajectory as the initial configuration for each of the three replicas. These random frames were verified to have an acyl chain RMSD of at least 5Å to generate substantial initial conformational diversity. We then run MT-REXEE simulations for the FabB-acyl complex in pocket A. As shown below, these simulations sample both pocket A and pocket B. We extract one frame in which each complex samples pocket B from these MT-REXEE simulations. We then ran 50 ns of standard MD on each of these pocket B configurations at $\lambda=0$ and again extracted a random frame from the first, second, and last third of the MD trajectories as the initial configurations.
@@ 3.2 Free Energy Protocol
We performed both EE and MT-REXEE simulations in triplicate for each system. The EE settings used were identical for both the MT-REXEE and independent EE simulations with a Wang-Landau scale of 0.82 and a Wang-Landau ratio of 0.85 for the equilibration for the FabB system, and 0.8 was used for both settings to calculate the free energy of ligand binding to MUP1. These parameters were tuned to ensure relatively even sampling between $\lambda$ intermediate states following weight equilibration. Slower weight equilibration through the increasing of Wang-Landau scale and ratio parameters extends equilibration times but produces more even $\lambda$ state sampling once equilibrated weights are obtained. The weight equilibration phase was ended when the Wang-Landau $\delta$ was less than 0.001 for both systems. The MUP1 ligands were run for 30 ns and 50 ns after equilibration for the solvent and protein complex simulations, respectively. For the FabB-acyl complexes, which represent an intermediate state preceded by FabB bound to acyl-ACP, we defined the“solvent” (i.e., pre-complex) simulation state as acyl-ACP in solvent. The FabB systems were run for 50 ns and 100 ns for the solvent and enzyme-bound states, respectively. All free energy estimates were calculated using the MBAR implementation in alchemlyb 42.
We also ensured that the overlap between adjacent $\lambda$ states was sufficient for all transformations both in solvent (Figure S8) and in complex (Figure S9). We used the same $\lambda$ state spacing for both the EE and MT-REXEE simulations, as well as the same initial configurations for all replicas.
The primary feature added to the MT-REXEE package in this study is the use of a new swapping criterion between simulations. In the previously developed exhaustive approach, swaps are only attempted for the last frame of each iteration. This new swapping criterion is named random range (RR), and rather than only allowing swaps between trajectories in the last frame of an iteration, we instead allow swaps at random compatible frames (i.e. endpoints of two neighboring simulations that represent, up to the presence of different dummy atoms, the same physical system) in a defined portion of the trajectory. We used the last 50% of the trajectory to result in longer continuous sampling within each trajectory, decreasing the structural correlation to the previous iteration. Regardless of the location of the trajectory used for swapping, all data from each iteration is used to compute free energy estimates and any additional analysis; only trajectory swapping and continuation use the randomly sampled exchange-compatible frame. For the MUP1 system, we used redundant end states (5 intermediate $\lambda$ states as well as 5 states each at $\lambda$=0 and $\lambda$=1) in combination with the exhaustive swapping method as we did in our previous MT-REXEE paper 14 to compare to the random range swapping approach without redundant end states. For the FabB-acyl systems, we ran only MT-REXEE simulations with the RR method and with single $\lambda$ end-states.
@@ 3.3 Configurational Analysis
X-ray crystal structures of the FabB–acyl-ACP complex show two distinct binding pockets, for the acyl chain. We speculated that these two pockets, termed A and B, would also be occupied in the covalent intermediate. To develop an unbiased classifier of pocket occupancy, we used the MT-REXEE simulations with the acyl chain initialized in the canonical binding pocket (A) and visually labeled a subset of trajectory frames (12–15) for all chain lengths as being in either pocket A or B. We determined that the distance between residue THR300 and carbon 4 in the acyl chain could be used to differentiate the two pockets (Figure S3A). We evaluated several distance thresholds between pairs of atoms in the acyl chain and FabB as well as acyl internal dihedrals as metrics for pocket determination, but the distance between C4 and THR300 had the least overlap between pockets A and B. We plotted the distance distribution of these residues in EE simulations, which were visually verified to have all frames within either pocket A or B, as well as for frames from our MT-REXEE simulations, which sample both pockets, to determine the threshold of 0.85 nm (Figure S3B). Several conformations for longer chains (8 or more carbons) were identified as belonging in pocket A or B by this criterion, but did not match existing known structures for the pockets. Thus, we also applied a chain compaction criterion that the distance between carbons 2 and 8 should be shorter than 0.6 nm in pocket B and longer than 0.75 nm in pocket A, limits based on crystal structures. Any trajectory frames in which this distance to THR300 is greater than 0.85 nm and the compaction distance is greater than 0.75 nm are considered in pocket A. Any frames with a distance to THR300 less than 0.85 nm and a compaction distance of less than 0.6 nm are considered in pocket B. Any frames for which neither pocket criteria are satisfied were discarded from free energy analysis, as they are in neither pocket. For the C2 and C4 complexes, these discarded frames were 1-4% of the trajectories, and for all other chain lengths, the discarded frames were << 0.5% of the trajectories.
@@ 4 Results
4.1 Random range swapping improves round-trip time and statistical convergence compared to previous exchange methods
@@
In this section, we compare the results for running binding free energy calculations for a set of small molecule ligands with the protein MUP1 using both the previously introduced exhaustive method and the random range (RR) method (Figure 2A–B). We ran this comparison to determine any changes in performance and ensure that this new algorithm for selecting swappable states does not introduce additional error. The RR method no longer requires redundant end states. This eliminates the increase in Wang-Landau weight equilibration times observed in our earlier work 14. The result of this improvement is that MT-REXEE with the RR exchange method (MT-REXEE RR) has equivalent or lower total Wang-Landau weight equilibration times compared to independent EE simulations (Figure S4). We note that weight equilibration times are calculated using the full iteration in which the weights reach the equilibration criteria, even if conformational exchanges occur at a frame earlier than the final frame of the iteration. Only the E-F transformation had a potentially statistically significant difference between the weight equilibration times using EE and MT-REXEE RR ($p=0.083$). We hypothesize that the decrease in weight equilibration times for the E-F transformation is likely due to the improved conformational space sampling using RR. This transformation is the most chemically distinct of all the MUP1 ligands evaluated in this study, and for this reason, it benefits most from the increased conformational sampling.
@@
Figure 2: Random range swapping yields statistically equivalent free energies while significantly decreasing the round-trip time through alchemical states. (A) A diagram of the sequence of ligand transformations used in binding free energy estimates with the protein MUP1 to validate the random range swapping scheme. The curved arrows show where MT-REXEE swaps occur. (B) The structure of MUP1 in complex with ligand A. (C) We compare the binding free energy estimates using $E$, MT-REXEE with the exhaustive swapping scheme, and MT-REXEE with the RR swapping scheme. We observe no significant differences between the free energy estimates for all three methods as the values are all within statistical error of one another. (D) We observe a significant (p<<0.001) decrease in the round-trip times from 10-20 ns to 3-5 ns for both solvent and MUP1 complex simulations using the RR swapping scheme compared to the previous exhaustive scheme.
@@
The binding free energy estimates are statistically equivalent for EE and MT-REXEE with both exhaustive and RR swapping methods, but various convergence metrics indicate that less simulation time is necessary for equal accuracy with RR (Figure 2C). We measured the convergence of free energy estimates both along the length of individual trajectories as well as convergence between trajectories initiated with significantly different initial configurations. Importantly, we observe that for both solvent and, even more significantly for the protein complex simulations, the free energy estimates converge significantly more quickly, resulting in closer estimates between the estimates for the first and last half of the trajectories for MT-REXEE RR compared to both EE and the MT-REXEE simulations using the exhaustive method (Figure S5A–B). We also notice a significant decrease in the standard error of the estimate between replicates of between 40 and 160% compared to EE (Figure S5C). These independent replicates were initiated with the same initial configurations extracted from standard MD trajectories separated by a minimum of 20 ns. The exchanges between distinct alchemical transformations in the MT-REXEE method enable increased sampling of the conformational space, which produces faster convergence. The standard error is also reduced by between 16-90% when using the exhaustive swapping method with redundant end states compared to EE. This is likely because the increase in conformational sampling is balanced by the decreased sampling of intermediate $\lambda$ states due to the addition of redundant end states. This results in decreased conformational sampling than MT-REXEE RR, but still faster conformational sampling than EE. This is consistent with the observed significant decrease in the round-trip times for the MT-REXEE RR method compared to the exhaustive swapping method (Figure 2D). We expect the over 4-fold decrease in round-trip time is primarily responsible for the faster increased conformational sampling in MT-REXEE RR compared to the MT-REXEE exhaustive method.
@@ 4.2 Enhanced Sampling of FabB with MT-REXEE
We apply the MT-REXEE method to an example system in which there is a known sampling limitation. Prior evidence suggests that when FabB binds acyl-ACP substrates, the acyl chain samples two distinct binding tunnels, pockets A and pocket B, in addition to a transient binding site, pocket C, where the acyl chain is parallel to the Ppant arm–primarily in the enzyme-ACP complex. Here we examine the FabB-acyl intermediate–that is, the complex formed after transfer of the acyl chain from the phosphopantetheine of ACP to the catalytic CYS163 of FabB (Figure 3A). Sampling the FabB-acyl complex is simpler than sampling the FabB–acyl-ACP complex, as the motions of ACP produce additional sampling issues; the FabB–ACP interface is large and flexible. Using the FabB-acyl complex allows us to focus on only the motions of the acyl chain. Additionally, the FabB-acyl complex precedes the condensation reaction with malonyl-ACP, which extends the acyl chain by two carbons, and is a biochemically important state 8.
@@
Previous mutational data of FabB suggest that pocket B remains accessible in the covalent intermediate, as mutants that block binding to pocket A are still capable of elongating acyl chains up to 16 carbons in length 24. The transition between pockets A and B is too slow to sample with direct simulation using standard MD with reasonable simulation lengths (500 ns; Figure 3B). Therefore, this study uses MT-REXEE with random range swapping to evaluate the relative stability of 2, 4, 6, 8, 10, 12, 14, and 16 carbon length acyl chain substrates covalently bound to FabB. We simulated FabB as a dimer with only one substrate attached. This approach is motivated by previous work on multimers of FAS enzymes, where binding of multiple substrates can yield negative cooperativity between monomers 25, 33.
@@
We performed the 500 ns standard MD simulations for FabB-acyl complexes with acyl substrates from 2–14 carbons in length, seven simulations in all. All simulations were initiated with the same initial configuration as the corresponding MT-REXEE simulation in pocket A, with the removal of dummy atoms required for MT-REXEE. We ran three replicates with distinct initial configurations for all chain lengths (see Methods). The corresponding MT-REXEE simulations were run for only 100 ns (in parallel) for each transformation of chain length. We then analyzed the trajectories to determine the occupancy of pocket A, B, or neither for each chain length substrate. For MT-REXEE simulations, we specifically compared the pocket occupancy for the end states and not for intermediate $\lambda$ states, as these are irrelevant to the physical occupancies.
@@
Figure 3: (A) The binding pocket of the FabB–acyl covalent intermediate is shown with a 12 carbon acyl substrate in both pocket A (orange) and pocket B (brown) in the context of the overall protein. The structures shown are the centroids from MT-REXEE simulations of the C10-12 transformation for $\lambda=1$. (B) Occupancy of pocket B as a function of substrate chain length using both standard MD simulations and MT-REXEE enhanced sampling as described in section $3.$ Standard MD is not able to access pocket B within 500 ns for any substrate longer than 4 carbons, but we readily see sampling in pocket B for all chain lengths within only 100 ns of MT-REXEE simulation. (C) A close-up image of the binding pocket in (A) highlighting several residues on both the catalytically active monomer and adjacent monomer that restrict the acyl chain’s conformational flexibility and prevent interconversion during MD simulations of at least 500 ns.
@@
MT-REXEE and standard MD show statistically equivalent sampling of pocket B only for the 2 carbon acyl chain substrate. For all longer substrates, standard MD shows reduced sampling of pocket B, likely because of the high free energy barrier for longer substrates to transition to this pocket (Figure 3B). In contrast, MT-REXEE enables significant sampling in pocket B for all chain lengths studied while utilizing only 20% of the trajectory length as the standard MD simulations.
@@
We hypothesize that the inability of standard MD to sample both pockets results from stabilizing interactions between the protein and acyl chain that are present for chains longer than 4 carbons (Figure 3C). For both MD and MT-REXEE simulations, the 2-carbon acyl chains sample pockets A and B with relatively high frequency. The presence of the conformational swaps between the 2 carbon state and longer chain substrates allows for conformational sampling of both pockets at all chain lengths. MT-REXEE enables simulation trajectories to sample pocket B when simulations are initiated in pocket A, thus requiring no prior knowledge of the presence or location of the alternate pocket. In contrast, existing knowledge of both the location of pocket B and a reasonable transition path would be necessary for employing a configurational variable dependent method. The lack of directly applied configurational bias along a known CV, like that used in metadynamics, also provides increased confidence that these alternate conformations are not artifacts of the enhanced sampling method itself.
@@
In MT-REXEE simulations, switching between pockets is primarily limited by the round-trip time between alchemical simulations of different chain lengths, as we can only switch between pockets in simulations of the shortest chain. Additionally, we want to emphasize that the success of the MT-REXEE method is heavily correlated to the suitability of the applied equilibrated EE weights. An increase in round-trip time correlates with an increased imbalance between $\lambda$ state sampling within the simulations (Figure S6). The imbalance in $\lambda$ state sampling can be minimized by adjusting Wang-Landau equilibration parameters. Optimization of these parameters is system-specific. Further optimization of these parameters would likely further minimize this effect, but would also increase weight equilibration times. This necessitates a balance between the efficiency of weight equilibration and the accuracy of equilibrated weights.